Please Answer a-dQuestion Help In the last quarter of 2007, a group of 64 mutual funds had a mean return of 2.5% with a standard deviation of
Answer #1

a)

probability = P(X>14.8) = P(Z>3)= 0.15 %

b)

probability = P(X<2.5) = P(Z<0)= 50.0 %

c)

probability = P(-1.6<X<6.6) = P(-1<Z<1)= 68 %

d)

probability = P(X>10.7) = P(Z>2)= 2.5 %